# Lyapunov exponents and geodesic stability of Schwarzschild black hole in the non-commutative gauge theory of gravity

### v3

@@ Abstract
In this paper, we study the stability of geodesic motion for both massive and massless particles using Lyapunov exponents in the non-commutative (NC) Schwarzschild black hole (BH) via the gauge theory of gravity. As a first step, we investigate the both time-like and null radial motion of particles, the mean result in NC geometry shows that the particles take infinity proper time to reach the NC singularity (infinite time affine parameter framework for photons). The proper/coordinate time of Lyapunov exponents and their ratio of time-like geodesic for the circular motion of this black hole shows a new behavior, which describes a new range of stable circular orbits between unstable ones. Then we analyze the circular motion of photons, where the result shows a new photon sphere near the event horizon which is not allowed in the commutative case, and the Lyapunov exponent is expressed in this geometry, where this confirms the instability of the outer photon sphere and the stability of the inner one. Moreover, we studied the effect of noncommutativity on the black hole shadow radius, We found a similarity between the non-commutativity and the mass of a black hole. Then we using experimental data from the event horizon telescope, we show that a noncommutativity parameter of the order of $\Theta^{\text{Phy}}\sim 10^{-32}m$.
@@ I Introduction
The investigation of geodesic motion for test particles around compact celestial bodies is fundamental in astrophysics, offering valuable insights into the nature of neutron stars, black holes (BHs), and other dense astrophysical objects. The trajectory of particles moving around a BH encodes information about the underlying spacetime geometry, which is shaped by the strong gravitational field of the compact object. This study focuses on analyzing how non-commutative (NC) geometry influences the geodesic motion of test particles near a Schwarzschild BH, employing the Lyapunov exponent Lyapunov5 as a diagnostic tool. Two key types of geodesic motion often examined in BH physics are radial and circular trajectories. A wide range of studies has previously explored geodesic motion around BHs, such as Chandr1 ; gibbons1 ; islam1 ; jaklit ; stuchli ; kran1 ; kran2 ; cruz ; hackmann1 ; kagraman ; hackmann2 ; hackmann3 ; levin ; grunau ; belbrun ; barack ; pugliese1 ; pugliese2 ; virb1 ; virb2 ; pradhan ; olivares ; villanueva ; pugl1 ; pugl2 . The classification of stable and unstable orbits sheds light on the spacetime geometry surrounding these objects. The Lyapunov exponent $\lambda$ provides a means to evaluate the stability of equatorial circular geodesics in BH spacetimes. Depending on the nature of the orbits—whether stable, unstable, or marginally stable—the Lyapunov exponent helps in determining their properties. Circular geodesics in a Schwarzschild BH are integrable, meaning they exhibit no chaotic behavior. In particular, unstable orbits have a positive Lyapunov exponent Lyapunov5 , while stable and marginally stable orbits correspond to imaginary and zero values of the exponent, respectively Lyapunov1 ; Lyapunov2 ; Lyapunov3 ; Lyapunov5 ; Lyapunov6 . The importance of geodesic stability has driven numerous studies that utilize Lyapunov exponents to explore various BH spacetimes geostab12 ; geostab11 ; geostab3 ; geostab8 ; geostab9 ; geostab4 ; geostab10 ; geostab6 ; geostab1 ; geostab2 ; geostab7 .
@@
Recent research has investigated how non-commutativity alters geodesic dynamics using multiple approaches, such as modeling the NC matter distribution through Gaussian functions nozari1 ; nozari2 ; kuniyal ; iftikhar ; larranaga ; bhar ; rahaman and Lorentzian profiles NCLD1 ; NCLD2 ; NCLD3 , in addition to the star product formalism as well as Bopp’s transformation ulh1 ; larbi1 and NC gauge gravity abdellah1 ; abdellah4 ; NCGTG2 ; NCGTG3 ; NCGTGmassdeformed1 ; NCGTGmassdeformed2 ; NCGTGmassdeformed3 . In previous work abdellah1 , we identified a new class of particle motion near the event horizon in an NC spacetime for massive test particles. Here, we extend this analysis to massless particles and other motion types, providing a more comprehensive study of geodesic stability. This work contributes to the ongoing exploration of orbital stability using Lyapunov exponents in NC spacetimes geostab14 ; geostab13 ; geostab5 .
@@
The coordinate operators in NC spacetime obey the following commutation relation:
@@
where $\Theta^{\mu\nu}$ represents an antisymmetric real matrix that sets the fundamental discretization scale of spacetime. This modification affects the Heisenberg uncertainty principle, leading to a minimum measurable length on the order of the Planck scale:
@@
In the NC framework, the usual function multiplication is replaced by the Moyal star product ”*”, which is defined for two functions $f(x)$ and $g(x)$ as follows:
@@
In this study, we adopt the NC gauge gravity approach in conjunction with the Seiberg-Witten (SW) map seiberg1 and star products. NC gauge gravity is a candidate extension of general relativity, maintaining action invariance under NC transformations via the SW map action1 ; action2 ; action3 . We utilize this framework to derive the NC Schwarzschild metric, then compute the NC corrections to the effective potential and analyze geodesic motion for different particle types. Our findings suggest that both massive and massless particles are prevented from reaching the NC singularity within a finite proper time or affine parameter, while also revealing the presence of novel stable circular orbits near the event horizon—distinct from those predicted in other NC models. Further, we assess the stability of these orbits through Lyapunov exponent analysis, confirming their stability in NC spacetime. Additionally, we determine NC corrections to the BH shadow radius and use observational data to estimate the NC parameter, yielding a fundamental length close to the Planck scale.
@@
The structure of this paper is as follows: Section II details the NC corrections to the metric field using the star product and SW maps. Section III derives the NC geodesic equations and the effective potential, incorporating second-order corrections in $\Theta$. We examine radial and circular motion for both massive and massless particles, studying orbital stability through the Lyapunov exponent. This is followed by an exploration of NC corrections to the BH shadow radius and an estimation of the NC parameter using observational data. The final section summarizes our key results and conclusions.
II Non-commutative corrections for Schwarzschild black hole
@@
In this section, we briefly present the NC correction to the Schwarzschild metric within the framework of the gauge theory of gravity. Following the methodology outlined in our previous works abdellah1 ; abdellah6 ; abdellah2 , we employ both the Seiberg-Witten (SW) map and the star ($\ast$)-product to construct the NC metric for the Schwarzschild black hole. To obtain the NC gauge potential of gravity, represented by the tetrad fields, we utilize the perturbative expansion of the SW map for the tetrad fields $\hat{e}^{a}_{\mu}(x,\Theta)$. This expansion is carried out as a power series in $\Theta$ up to second order, as derived in Ref. cham1 .
$\hat{e}_{\mu}^{a}=\,$ $e_{\mu}^{a}-\frac{i}{4}\Theta^{\nu\rho}[\omega^{ac}_{\nu}\partial_{\rho}e^{d}_{\mu}+(\partial_{\rho}\omega^{ac}_{\mu}+R^{ac}_{\rho\mu})e^{d}_{\nu}]\eta_{cd}+\frac{1}{32}\Theta^{\nu\rho}\Theta^{\lambda\tau}\left[2\{R_{\tau\nu},R_{\mu\rho}\}^{ab}e^{c}_{\lambda}-\omega^{ab}_{\lambda}(D_{\rho}R_{\tau\mu}^{cd}+\partial_{\rho}R_{\tau\mu}^{cd})e^{m}_{\nu}\eta_{dm}\right.$
$\left.-\{\omega_{\nu},(D_{\rho}R_{\tau\mu}+\partial_{\rho}R_{\tau\mu})\}^{ab}e^{c}_{\lambda}-\partial_{\tau}\{\omega_{\nu},(\partial_{\rho}\omega_{\mu}+R_{\rho\mu})\}^{ab}e^{c}_{\lambda}-\omega^{ab}_{\lambda}\left(\omega^{cd}_{\nu}\partial_{\rho}e^{m}_{\mu}+\left(\partial_{\rho}\omega_{\mu}^{cd}+R_{\rho\mu}^{cd}\right)e^{m}_{\nu}\right)\eta_{dm}\right.$
$\left.+2\partial_{\nu}\omega_{\lambda}^{ab}\partial_{\rho}\partial_{\tau}e^{c}_{\mu}-2\partial_{\rho}\left(\partial_{\tau}\omega_{\mu}^{ab}+R_{\tau\mu}^{ab}\right)\partial_{\nu}e^{c}_{\lambda}-\{\omega_{\nu},(\partial_{\rho}\omega_{\lambda}+R_{\rho\lambda})\}^{ab}\partial_{\tau}e^{c}_{\mu}-\left(\partial_{\tau}\omega_{\mu}^{ab}+R^{ab}_{\tau\mu}\right)\left(\omega^{cd}_{\nu}\partial_{\rho}e^{m}_{\lambda}\right.\right.$
$\left.\left.+\left((\partial_{\rho}\omega^{cd}_{\lambda}+R^{cd}_{\rho\lambda})\right)e^{m}_{\nu}\right)\eta_{dm}\right]\eta_{cb},$ (5)
and the commutative gauge gravity potentials are denoted by $\hat{e}_{a}^{\mu}$ and $\omega^{ab}_{\mu}$, which are the tetrad fields and the spin connection, and:
@@
$D_{\mu}R_{\rho\sigma}^{ab}=\partial_{\mu}R^{ab}_{\rho\sigma}+$ $\left(\omega_{\mu}^{ac}R^{db}_{\rho\sigma}+\omega_{\mu}^{bc}R^{da}_{\rho\sigma}\right)\eta_{cd}.$ (7)
The component $\hat{e}_{a}^{\mu}$ is the of the vierbein $\hat{e}_{\mu}^{a}$ defined as:
$\hat{e}_{\mu}^{b}\hat{e}_{a}^{\mu}=\delta_{a}^{b},\quad\hat{e}_{\mu}^{a}\hat{e}_{a}^{\nu}=\delta_{\mu}^{\nu}\,.$ (8)
In this context, the NC correction to the metric component $\hat{g}_{\mu\nu}$ is obtained according to the following definition chai1 :
$\hat{g}_{\mu\nu}=\frac{1}{2}(\hat{e}_{\mu}^{b}\ast\hat{e}_{\nu b}+\hat{e}_{\nu}^{b}\ast\hat{e}_{\mu b})\,.$ (9)
In this work we choose the following NC matrix $\Theta^{\mu\nu}$:
@@
In order to obtain the NC correction to the metric we follow the same steps in Ref. abdellah1 , and we take the same choose of the general tetrads field:
$\underline{e}_{\mu}^{0}$ $=\left(\begin{array}[]{cccc}\left(1-\frac{2m}{r}\right)^{\frac{1}{2}},&0,&0,&0\end{array}\right),$ (11b)
$\underline{e}_{\mu}^{1}$ $=\left(\begin{array}[]{cccc}0,&\left(1-\frac{2m}{r}\right)^{-\frac{1}{2}}\sin\theta\cos\phi,&r\,\cos\theta\cos\phi,&-r\,\sin\theta\sin\phi\end{array}\right),$ (11d)
$\underline{e}_{\mu}^{2}$ $=\left(\begin{array}[]{cccc}0,&\left(1-\frac{2m}{r}\right)^{-\frac{1}{2}}\sin\theta\sin\phi,&r\,\cos\theta\sin\phi,&r\,\sin\theta\cos\phi\end{array}\right),$ (11f)
$\underline{e}_{\mu}^{3}$ $=\left(\begin{array}[]{cccc}0,&\left(1-\frac{2m}{r}\right)^{-\frac{1}{2}}\cos\theta,&-r\,\sin\theta,&0\end{array}\right).$ (11h)
The NC tetrads $\hat{e}^{a}_{\mu}$ are abdellah1 :
$\hat{e}^{0}_{0}=$ $\left(1-\frac{2m}{r}\right)^{\frac{1}{2}}+\Theta^{2}\Bigg{\{}\frac{m\left(88m^{2}+mr\left(-77+15\sqrt{1-\frac{2m}{r}}\right)-8r^{2}\left(-2+\sqrt{1-\frac{2m}{r}}\right)\right)\sin^{2}\theta}{32r^{5}(1-\frac{2m}{r})^{3/2}}\Bigg{\}}+\mathcal{O}(\Theta^{3})$ (12a)
$\hat{e}^{1}_{1}=$ $\left(1-\frac{2m}{r}\right)^{-\frac{1}{2}}sin\theta cos\phi-\frac{im\,\Theta}{4r^{2}}\left(1-\frac{2m}{r}\right)^{-\frac{3}{2}}sin\theta sin\phi+\Theta^{2}\Bigg{\{}\frac{m\Bigg{(}m\bigg{(}25+2\sqrt{1-\frac{2m}{r}}\bigg{)}r-2\bigg{(}1+\sqrt{1-\frac{2m}{r}}\bigg{)}r^{2}}{64r^{5}\big{(}1+\frac{2m}{r}\big{)}^{5/2}}$
$+\frac{-24m^{2}+\bigg{(}24m^{2}+m\big{(}-29+6\sqrt{1-\frac{2m}{r}}\big{)}\bigg{)}r+2\bigg{(}5-3\sqrt{1-\frac{2m}{r}})r^{2}\bigg{)}\bigg{)}\cos(2\theta)\Bigg{)}}{64r^{5}\big{(}1+\frac{2m}{r}\big{)}^{5/2}}\Bigg{\}}sin\theta cos\phi+\mathcal{O}(\Theta^{3})$ (12b)
$\hat{e}^{2}_{1}=$ $\left(1-\frac{2m}{r}\right)^{-\frac{1}{2}}sin\theta sin\phi+\frac{im\,\Theta}{4r^{2}}\left(1-\frac{2m}{r}\right)^{-\frac{3}{2}}sin\theta cos\phi+\Theta^{2}\Bigg{\{}\frac{m\Bigg{(}m\big{(}25+2\sqrt{1-\frac{2m}{r}}\big{)}r-2\bigg{(}1+\sqrt{1-\frac{2m}{r}}\bigg{)}r^{2}}{64r^{5}\big{(}1+\frac{2m}{r}\big{)}^{5/2}}$
$+\frac{-24m^{2}+\bigg{(}24m^{2}+m\bigg{(}-29+6\sqrt{1-\frac{2m}{r}}\bigg{)}\bigg{)}r+2\bigg{(}5-3\sqrt{1-\frac{2m}{r}})r^{2}\bigg{)}\bigg{)}\cos(2\theta)\Bigg{)}}{64r^{5}\big{(}1+\frac{2m}{r}\big{)}^{5/2}}\Bigg{\}}sin\theta sin\phi+\mathcal{O}(\Theta^{3})$ (12c)
$\hat{e}^{3}_{1}$ $=\left(1-\frac{2m}{r}\right)^{-\frac{1}{2}}cos\theta+\Theta^{2}\Bigg{\{}\frac{m\Bigg{(}-24m^{2}+\bigg{(}29-6\sqrt{1-\frac{2m}{r}}\bigg{)}r-2\bigg{(}5-3\sqrt{1-\frac{2m}{r}}\bigg{)}r^{2}\Bigg{)}}{32r^{5}\big{(}1+\frac{2m}{r}\big{)}^{5/2}}\Bigg{\}}sin^{2}\theta\cos\theta+\mathcal{O}(\Theta^{3})$ (12d)
$\hat{e}^{1}_{2}=$ $r\cos\theta\cos\phi+\frac{i\Theta}{4}\Bigg{(}1-\bigg{(}1-\frac{2m}{r}\bigg{)}^{-\frac{1}{2}}\Bigg{)}\cos\theta\sin\phi+\Theta^{2}\Bigg{\{}\frac{m\Big{(}\Big{(}6-9\sqrt{1-\frac{2m}{r}}\Big{)}r+\Big{(}-2+5\sqrt{1-\frac{2m}{r}}\Big{)}r\cos(2\theta)}{64r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}$
$+\frac{\bigg{(}18r+\bigg{(}-28+8\sqrt{1-\frac{2m}{r}}\bigg{)}m\bigg{)}\sin^{2}\theta}{64r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}\Bigg{\}}\cos\theta\cos\phi+\mathcal{O}(\Theta^{3})$ (12e)
$\hat{e}^{2}_{2}=$ $r\cos\theta\sin\phi-\frac{i\Theta}{4}\Bigg{(}1-\bigg{(}1-\frac{2m}{r}\bigg{)}^{-\frac{1}{2}}\Bigg{)}\cos\theta\cos\phi+\Theta^{2}\Bigg{\{}\frac{m\Big{(}\Big{(}6-9\sqrt{1-\frac{2m}{r}}\Big{)}r+\Big{(}-2+5\sqrt{1-\frac{2m}{r}}\Big{)}r\cos(2\theta)}{64r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}$
$+\frac{\bigg{(}18r+\bigg{(}-28+8\sqrt{1-\frac{2m}{r}}\bigg{)}m\bigg{)}\sin^{2}\theta}{64r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}\Bigg{\}}cos\theta sin\phi+\mathcal{O}(\Theta^{3})$ (12f)
$\hat{e}^{3}_{2}=$ $-r\sin\theta+\Theta^{2}\Bigg{\{}\frac{m\bigg{(}-2m\bigg{(}1+2\sqrt{1-\frac{2m}{r}}\bigg{)}+\bigg{(}m\bigg{(}-14+4\sqrt{1-\frac{2m}{r}}\bigg{)}+\bigg{(}11-5\sqrt{1-\frac{2m}{r}}\bigg{)}r\bigg{)}\cos(2\theta)}{64r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}$
$+\frac{r\bigg{(}1+\sqrt{1-\frac{2m}{r}}\bigg{)}\bigg{)}}{64r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}\Bigg{\}}\sin\theta+\mathcal{O}(\Theta^{3})$ (12g)
$\hat{e}^{1}_{3}=$ $-rsin\theta sin\phi+\frac{i\Theta}{4}\Bigg{(}1-\bigg{(}1-\frac{2m}{r}\bigg{)}^{-\frac{1}{2}}\Bigg{)}\sin\theta\cos\phi+\Theta^{2}\Bigg{\{}\frac{m\bigg{(}-2r\bigg{(}1-\sqrt{1-\frac{2m}{r}}\bigg{)}+\bigg{(}2m\bigg{(}9-17\sqrt{1-\frac{2m}{r}}\bigg{)}\bigg{)}}{32r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}$
$+\frac{2r-\bigg{(}13-17\sqrt{1-\frac{2m}{r}}\bigg{)}r\bigg{)}\sin^{2}\theta}{32r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}\Bigg{\}}\sin\theta\sin\phi+\mathcal{O}(\Theta^{3})$ (12h)
$\hat{e}^{2}_{3}=$ $rsin\theta cos\phi+\frac{i\Theta}{4}\Bigg{(}1-\bigg{(}1-\frac{2m}{r}\bigg{)}^{-\frac{1}{2}}\Bigg{)}\sin\theta\sin\phi-\Theta^{2}\Bigg{\{}\frac{m\bigg{(}-2r\bigg{(}1-\sqrt{1-\frac{2m}{r}}\bigg{)}+\bigg{(}2m\bigg{(}9-17\sqrt{1-\frac{2m}{r}}\bigg{)}\bigg{)}}{32r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}$
$+\frac{2r-\bigg{(}13-17\sqrt{1-\frac{2m}{r}}\bigg{)}r\bigg{)}\sin^{2}\theta}{32r^{3}\big{(}1+\frac{2m}{r}\big{)}^{3/2}}\Bigg{\}}\sin\theta\cos\phi+\mathcal{O}(\Theta^{3}).$ (12i)
Noting that, in the limit of $\Theta=0$, we recover the commutative tetrad fields (11b-11h).

### v4

@@ Abstract
In this paper, we study the stability of geodesic motion for both massive and massless particles using Lyapunov exponents in the non-commutative (NC) Schwarzschild black hole (BH) via the gauge theory of gravity. As a first step, we investigate both time-like and null radial motion of particles. The main result in NC geometry shows that the particles take an infinite proper time to reach the NC singularity (infinite affine parameter time framework for photons). The proper time and coordinate time of Lyapunov exponents, as well as their ratio for time-like geodesics in circular motion around this black hole, exhibit a new behavior, revealing a new range of stable circular orbits between unstable ones. Next, we analyze the circular motion of photons, where the results indicate the presence of a new photon sphere near the event horizon, which is not allowed in the commutative case. The Lyapunov exponent in this geometry confirms the instability of the outer photon sphere and the stability of the inner one. Moreover, we study the effect of non-commutativity on the black hole shadow radius and find a similarity between non-commutativity and the mass of a black hole. Finally, using experimental data from the Event Horizon Telescope, we show that the non-commutativity parameter is of the order of $\Theta^{\text{Phy}}\sim 10^{-34}\,m$.
@@ I Introduction
The investigation of geodesic motion for test particles around compact celestial bodies is fundamental in astrophysics, offering valuable insights into the nature of neutron stars, black holes (BHs), and other dense astrophysical objects. The trajectory of particles moving around a BH encodes information about the underlying spacetime geometry, which is shaped by the strong gravitational field of the compact object. This study focuses on analyzing how non-commutative (NC) geometry influences the geodesic motion of test particles near a Schwarzschild BH, employing the Lyapunov exponent Lyapunov5 as a diagnostic tool. Two key types of geodesic motion often examined in BH physics are radial and circular trajectories. A wide range of studies has previously explored geodesic motion around BHs pradhan ; villanueva ; pugl1 ; pugl2 ; Geodesic1.1 ; Geodesic3.1 ; Geodesic5.3 . The classification of stable and unstable orbits sheds light on the spacetime geometry surrounding these objects. The Lyapunov exponent $\lambda$ provides a means to evaluate the stability of equatorial circular geodesics in BH spacetimes. Depending on the nature of the orbits whether stable, unstable, or marginally stable the Lyapunov exponent helps in determining their properties. Circular geodesics in a Schwarzschild BH are integrable, meaning they exhibit no chaotic behavior. In particular, unstable orbits have a positive Lyapunov exponent Lyapunov5 , while stable and marginally stable orbits correspond to imaginary and zero values of the exponent, respectively Lyapunov1 ; Lyapunov5 . The importance of geodesic stability has driven numerous studies that utilize Lyapunov exponents to explore various BH spacetimes geostab12 ; geostab11 ; geostab3 ; geostab9 ; geostab4 ; geostab1 ; geostab2 ; geostab7 .
Another critical feature of geodesic motion around BH spacetimes is the photon sphere, a region where photons are forced into circular orbits due to the strong gravitational field. This region plays a significant role in understanding the behavior of light near compact objects and is closely tied to the study of unstable circular orbits. The photon sphere also has implications for observables such as the BH shadow ehc , which is influenced by the dynamics of photons in this region geostab3 ; photonsphere1 ; photonsphere2 ; photonsphere4 ; radial1 ; photonspheres ; BHShadow1 .
@@
Recent research has investigated how non-commutativity alters geodesic dynamics using multiple approaches, such as modeling the NC matter distribution through Gaussian functions kuniyal ; iftikhar ; rahaman ; NCGD1 ; NCGD3 and Lorentzian profiles NCLD2 ; NCLD3 ; NCLD1 , in addition to the star product formalism, as well as Bopp’s transformation ulh1 ; larbi1 and NC gauge gravity abdellah1 ; abdellah4 ; NCGTG2 ; NCGTG3 ; NCGTGmassdeformed1 ; NCGTGmassdeformed2 ; NCGTGmassdeformed3 . In previous work abdellah1 , we identified a new class of particle motion near the event horizon in an NC spacetime for massive test particles. Here, we extend this analysis to massless particles and other motion types, providing a more comprehensive study of geodesic stability. In addition, in Ref.Tajron1 , the authors showed the existence of a new correction for the tetrad fields in the gravitational gauge theory, which was not present in the old model cham1 . We have taken this new correction into account in our study. This work contributes to the ongoing exploration of orbital stability using Lyapunov exponents in NC spacetimes geostab14 ; geostab13 ; geostab5 .
@@
In non‑commutative geometry, the position operators obey the relation:
@@
where $\Theta^{\mu\nu}$ an antisymmetric matrix that establishes a fundamental discretization scale for spacetime. This modification affects the Heisenberg uncertainty principle, leading to a minimum measurable length on the order of the Planck scale:
@@
In this geometry, the usual function multiplication is replaced by the Moyal star product ”*”, which is defined for two functions $f(x)$ and $g(x)$ as follows:
@@
In this study, we adopt the NC gauge gravity approach in conjunction with the Seiberg-Witten (SW) map seiberg1 and star products. NC gauge gravity is a candidate extension of general relativity, maintaining action invariance under NC transformations via the SW map action1 ; action3 . We utilize this framework to derive the NC Schwarzschild metric, then compute the NC corrections to the effective potential and analyze geodesic motion for different particle types. Our findings suggest that both massive and massless particles are prevented from reaching the NC singularity within a finite proper time or affine parameter while also revealing the presence of novel stable circular orbits near the event horizon distinct from those predicted in other NC models. Furthermore, we assess the stability of these orbits through Lyapunov exponent analysis, confirming their stability in NC spacetime. Additionally, we determine NC corrections to the BH shadow radius and use observational data to estimate the NC parameter through different phenomena, yielding a fundamental length close to the Planck scale.
@@
The structure of this paper is as follows: Section II overview the NC corrections to the metric field using the star product and SW maps. Section III derives the NC correction to both the geodesic equations and the effective potential, incorporating second-order corrections in $\Theta$. We examine radial and circular motion for both massive and massless particles, studying orbital stability through the Lyapunov exponent. This is followed by an exploration of NC corrections to the BH shadow radius and studying the limit of the NC parameter using observational data for different phenomena. The final section summarizes our key results and conclusions.
II Non-commutative corrections for Schwarzschild metric
@@
To construct the NC metric for the Schwarzschild BH we follow the methodology outlined in our previous works abdellah1 ; abdellah6 ; abdellah2 , and employ both the Seiberg-Witten (SW) map and the star ($\ast$)-product. To obtain the NC gauge potential of gravity, represented by the tetrad fields, we utilize the perturbation expansion of the SW map for the tetrad fields $\hat{e}^{a}_{\mu}(x,\Theta)$. This expansion is carried out as a power series in $\Theta$ up to the second order, as derived in Ref. Tajron1 .
$\hat{e}^{a}_{\mu}(x,\Theta)=e^{a}_{\mu}(x)-i\Theta^{\nu\rho}e^{a}_{\mu\nu\rho}(x)+\Theta^{\nu\rho}\Theta^{\lambda\tau}e^{a}_{\mu\nu\rho\lambda\tau}(x)+\mathcal{O}(\Theta^{3})$ (5)
where the first and second order corrections are given respectively by
$e^{a}_{\mu\nu\rho}(x)=\,\frac{1}{4}[\omega^{ac}_{\nu}\partial_{\rho}e^{d}_{\mu}+(\partial_{\rho}\omega^{ac}_{\mu}+F^{ac}_{\rho\mu})e^{d}_{\nu}]\eta_{cd}$ (6)
and
$e^{a}_{\mu\nu\rho\lambda\tau}$ $=\frac{1}{16}\Bigg{[}\left(2\{F_{\tau\nu},F_{\mu\rho}\}^{ab}e^{c}_{\lambda}-\omega^{ab}_{\lambda}(D_{\rho}F_{\tau\mu}^{cd}+\partial_{\rho}F_{\tau\mu}^{cd})e^{m}_{\nu}\eta_{dm}-\{\omega_{\nu},(D_{\rho}F_{\tau\mu}+\partial_{\rho}F_{\tau\mu})\}^{ab}e^{c}_{\lambda}\right.$
$\left.-\partial_{\tau}\{\omega_{\nu},(\partial_{\rho}\omega_{\mu}+F_{\rho\mu})\}^{ab}e^{c}_{\lambda}-\omega^{ab}_{\lambda}\left(\omega^{cd}_{\nu}\partial_{\rho}e^{m}_{\mu}+\left(\partial_{\rho}\omega_{\mu}^{cd}+F_{\rho\mu}^{cd}\right)e^{m}_{\nu}\right)\eta_{dm}+2\partial_{\nu}\omega_{\lambda}^{ab}\partial_{\rho}\partial_{\tau}e^{c}_{\mu}\right.$
$\left.-2\partial_{\rho}\left(\partial_{\tau}\omega_{\mu}^{ab}+F_{\tau\mu}^{ab}\right)\partial_{\nu}e^{c}_{\lambda}-\{\omega_{\nu},(\partial_{\rho}\omega_{\lambda}+F_{\rho\lambda})\}^{ab}\partial_{\tau}e^{c}_{\mu}-\left(\partial_{\tau}\omega_{\mu}^{ab}+F^{ab}_{\tau\mu}\right)\left(\omega^{cd}_{\nu}\partial_{\rho}e^{m}_{\lambda}\right.\right.$
[…]
where
@@
$D_{\mu}F_{\rho\sigma}^{ab}=\partial_{\mu}F^{ab}_{\rho\sigma}+$ $\left(\omega_{\mu}^{ac}F^{db}_{\rho\sigma}-\omega_{\mu}^{db}F^{ac}_{\rho\sigma}\right)\eta_{cd}.$ (9)
where the commutative gauge gravity potentials are denoted by $e_{a}^{\mu}$ and $\omega^{ab}_{\mu}$, which are the tetrad fields and the spin connection respectively.
Then a symmetric real metric component $\hat{g}_{\mu\nu}$ can be constructed according to chai1 :
$\hat{g}_{\mu\nu}=\frac{1}{2}(\hat{e}_{\mu}^{a}\ast\hat{e}_{\nu a}^{\dagger}+\hat{e}_{\nu}^{a}\ast\hat{e}_{\mu a}^{\dagger})\,.$ (10)
where $\hat{e}_{\nu a}^{\dagger}$ represents the complex conjugate of the deformed tetrad fildes fields $\hat{e}_{\nu a}$.
In space-space non-commutativity, we choose the coordinate system so that the NC matrix $\Theta^{\mu\nu}$ has the following form
@@
In order to obtain the NC correction to the metric, we follow the same steps as in Ref. abdellah1 and adopt the same choice of the general tetrad fields:
[…]